OpenStudy (anonymous):

solve for the value of "A" when sin A=3.5 X and Cos A=5.5x

OpenStudy (anonymous):

tan A=sin A/cos A so tan A=3.5x/5.5x tan A=3.5/5.5 tan A=0.6363\[\tan^{-1} 0.6363\]=A next find z angle in table

OpenStudy (anonymous):

then the answer is 32.47 degrees

OpenStudy (anonymous):

thx u ermi